If Dr. Suzuki rides his bike to his office, he averages12 miles per hour. If he drives his car, heaverages 36 miles per hour. His time driving is 1/4 hour less than his time bicycling. How far is his office?
Answer by ankor@dixie-net.com(22740)

If he averages 12 miles per hour on his bike, and he averages 36 miles per hour
if he drives. His time driving is 1/4 hour less than his time bicycling. How far is his office?
:
Let t = time on the bike
Then (t-.25) = time in the car
:
It's easy to find t, then we can find the distance from that:
Write a distance equation; Distance = Speed * time
:
car dist = bike dist
36(t-.25) = 12t
36t - 9 = 12t
36t - 12t = +9
24t = 9
t = 9/24
t = .375 hrs
:
Find distance using time on the bike:
.375 * 12 = 4.5 miles
:
Check solution using time in the car:
(.375 -.25) * 36
.125 * 36 = 4.5 miles also
:
Make sense to you?